The delarray package^{∗}

David Carlisle carlisle@cs.man.ac.uk

1994/03/14

1 Examples

The addition to array.sty added in delarray.sty is a system of implicit \left \right pairs. If you want an array surrounded by parentheses, you can enter:

\begin{array}({cc}) . . .

( \begin{matrix}a & b \\ c & d \end{matrix} )

Similarly if an environment equivalent to Plain TEX’s \cases could be defined by:

\begin{array}\{{lL}. . . .

f(x) = \{ \begin{matrix}0 & if x = 0 \\ sin(x)/x & otherwise \end{matrix}

Here L is supposed to denote a column of left aligned L-R text. It may be defined

via: \newcolumntype{L}{>{$}l<{$}}, as discussed in array.sty. Note that as

the delimiters must always be used in pairs, the ‘.’ must be used to denote a ‘null delimiter’.

This feature is especially useful if the [t] or [b] arguments are also used. In these cases the result is not equivalent to surrounding the environment by \left… \right, as can be seen from the following example:

( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) not ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} )

 \begin{array}[t]({c}) 1\\2\\3 \end{array}
 \begin{array}[c]({c}) 1\\2\\3 \end{array}
 \begin{array}[b]({c}) 1\\2\\3 \end{array}
 \quad\mbox{not}\quad
 \left(\begin{array}[t]{c} 1\\2\\3 \end{array}\right)
 \left(\begin{array}[c]{c} 1\\2\\3 \end{array}\right)
 \left(\begin{array}[b]{c} 1\\2\\3 \end{array}\right)

∗This file has version number v1.01, last revised 1994/03/14.